IF tan(pie cosa)=cot(pie sina) prove that: cos(a-45)=1/2√2

IF  tan(pie cosa)=cot(pie sina)


prove that: cos(a-45)=1/2√2

Grade:11

3 Answers

Rahul askIITiansExpert.IITR
13 years ago
the problem is fairly simple

cot x = tan ( pie/2 - x ).

so, pie/2-piesina = piecosa

1/2-sina=cosa

cosa+sina=1/2

cosa/√2 +sina/√2 =1/2√2

cos(a-45)=1/2√2

hence

Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation.

DO APPROVE IF THE ANSWER HELPS YOU.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Regards,

Askiitians Experts

Rahul- IIT Roorkee
yash mittal
13 years ago

tan(piecosa) = cot(piesina)

yash mittal
13 years ago

tan(piecosa) = cot(piesina)                                                                                                                                              tan(piecosa) = tan(90-180sina)                                                                                                                                       cosa + sina= 90                                                                                                                                                             cos(45-a) =1by2root2  (multiplying and div both sides by 1by root2)                                                                                                                                                                                     hence proved